You deciding between purchasing an electric car and an all-gas model. The electric model sells for 38,000 and can drive 258 miles per full charge.The all−gas model gets an effective gas mileage of 22 MPG and sells for 38,000 and can drive 258 miles per full charge.The all−gas model gets an effective gas mileage of 22 MPG and sells for 22,000. Assuming you drive 15,000 miles a year, which of the two cars would be the more economical over ten years? I.e., calculate the cost of the car including purchase and ten years’ worth of driving. Assume: (1) an average gas price of $3.12 per gallon over the entire ten-year period (2) an average cost of electricity of 11.3 cents / kWh, and that one charge of the electric model requires 64 kWh.
Answer #1

Total distance covered in 10 years = 150000 x 10 miles = 150000 miles

Full charges required for 15000 mile drive = 150000/258 [since 1 full charge can drive 258 miles]

Cost of charges for 15000 miles= No. of full charges x electricity required per full chagre x price rate of electricity

=(150000/258) x 64 kWh x 11.3cents/kWh

= $4204.65

Total 10 year cost of electric car = $38000 + $4204.65 = $42204.65

Gas required for 150000 miles drive = 150000/22 [since milage = 22 MPG]

Cost of gas for 150000 miles = gallons of gas required x gas price per gallon

= (150000/22) x $3.12

= $21272.73

Total 10 year cost  of all-gas car = $22000 + $21272.73 = $43272.73

Therefore, the electric car is more economical over 10 years.